Home | Contact Us | Submit Troubleshooting Info | Search | Help


    Output | Software Solutions | Proofing | Color, Screening and Calibration | Scanning | Digital Photography | Consumables | POS


Thermal 2: Planned Maintenance Procedure

Strictly confidential. Do not leave at a customer site.

Planned maintenance (PM) is to be performed only by Creo service personnel or by trained and certified operators.

Time needed:  1.5 hours  (Device can be returned to customer within approximately 40 minutes)

External check | Imaging check | Internal check | Sample script results

Requirements:

Before you start

Print this document Click the print icon at the top of the page. Make sure you print it in landscape format.
Safety Before you start, make sure you are familiar with all device and laser safety concerns. For more information see Laser safety information.

External checks

External Checks Action Look For Procedure or

Parts

Perform a visual check. Visually inspect the outside of the head and the surrounding area. Check for any device-specific issues (for example, Quantum reservoir hose integrity, stretched or pinched cables, ablation dust on the head or around the drum, and so on)  
Check the coolant and the coolant reservoir.
  • Visually inspect the coolant.
  • Visually inspect the reservoir.
  1. Check that coolant levels are good and that there is no leakage.
  2. Ensure that the flow from the reservoir is adequate.
  3. Check the reservoir for signs of contamination and floating debris.
  4. Ensure that you are using the correct coolant (TH2/2.5 uses 60/40; TH2.7 uses 70/30)
  5. Check the color of the coolant. If it is discolored or has lost color, replace it with new coolant.

To replace the coolant, see the Chiller page

Coolant part numbers:

  • 60/40 Coolant (pink) PN 82-1032A
  • 70/30 Coolant (blue) PN 402-00056
Check dust filter on chiller or heat exchanger. Visually inspect the dust filter. Check for dust, debris, and damage. Clean or replace as necessary.
Check that the coolant connectors have been upgraded (FCO CA09695) Complete the procedure Replacing the Coolant Connectors on a TH2 head    
Check the O-ring on the coolant hose. Visually inspect the O-ring and hose.
  • Check that there is no damage on the connectors.
  • Replace the O-ring, if needed (part number 59-3766)
TH2.7 only. Replace the external chiller coolant filter. Replace coolant filter assembly 500-01665A on rear of chiller.
  • Check that there is no damage on the connectors.
  • Replace the O-ring, if needed (part number 59-3766)
Verify the processor setup. Ensure the processor meets operating specifications. Check the setup specifications using the appropriate documentation.

Imaging check

Note: If you are using a laptop, the results will be automatically stored there. You will need to copy them to the workstation.

Imaging Checks Action Look for Procedure or

Parts

Back up the NVS settings. In Service Shell, run: NVSDump.spt Make sure that the file is in the results folder. Note that location of the folder.  
Generate a reference plot.

Run the standard test plot (Creo Plate Test Target).

 

See the Creo Plate Test Target site.
Generate a reference stroke cal. Create a stroke cal graph. In Service Shell, run: BeamBalance.spt

  • Examine the BeamBalance.spt graph in Service Shell. This gives a baseline of the health of the head. Compare to previous graphs and keep for reference.
  • If BeamBalance.spt. fails, run BBSProfile.spt to check the laser profile. Troubleshoot as needed (refer to the Thermal 2 troubleshooting page).

Safety Warning: Follow Laser Safety Procedures during the next steps.

Check the laser current.
  1. Type the following commands:
    TH2> eng
    TH2> bpd mode on
    TH2> lvd mode local
    TH2> ldd mode calibrate ##.#
    Use the current wet point power.
    TH2> ldd info
  2. In the service log, record the laser current, laser on hours, and the head serial number.
  3. Type the following command:
    TH2> set ldd
  4. In the service log, record the nominal current.
  5. Type the following commands:
    TH2> ldd mode off
    TH2> ldd info

  6. Confirm that the laser is turned off.
  • Compare the laser current and nominal current values to previously recorded values.
  • If the laser current is greater than the nominal current, contact the Response Centre.
 
 

Remove the head from the device.
Place the head on a clean working surface.

Note: If you are working on a Magnus 800, you do not need to remove the head.

Turn off the device. All laser safety procedures must be followed.

Inspect and clean the debris removal nozzle. If the device is equipped debris removal, clean the debris removal nozzle.
See: Importance of lens cleaning and debris removal
  • Inspect the debris removal hoses and ensure that there isn’t any cracking or wear.
  • Make sure the pressure tap in the debris nozzle is clear of debris. This ensures that the vacuum sensor will work properly.

Note: Except for Magnus 800, it is not possible to clean the debris removal nozzle if the head is on the device.

Thermal 2 Debris Removal Nozzle Cleaning Procedure
Inspect and clean the final window.

Clean the final window and surrounding area. You will need the lens cleaning kit (Lens cleaning kit: 04-0012B BOM).

Note that you must replace the head if the window is scratched, cracked, or if the coating is coming off.
See: Importance of lens cleaning and debris removal

  • Leaving streaks can seriously impact imaging quality.
  • The window cannot be replaced in the field.

Note: Except for Magnus 800, it is not possible to clean the final window if the head is on the device.

  Install the head.
  1. Re-attach the debris removal nozzle.
  2. Re-install the head using the head installation procedure.
  3. Run a surface depth series and an exposure series. If you replaced the window, you must also run slope and curve series.
  4. Load a plate.
  5. Run EdgeMCE.spt (or EdgeMPE.spt for NEWS devices with MPE) to check edge detection.
  6. Run the Creo Plate Test Target. If there are problems with the image quality, refer to the appropriate media technical note.
Check the head installation.
  1. Enable the write laser.
  2. Load a plate.
  3. Type the following commands:
    Note: This applies to MCE versions built using master code dated 2002:April:18 or later. To check the master code version type >> list version and check the baseline date.
MCE MPE
>> access expert
>> plot headtest
>> head install on
>> plot stop
>> access field
>> drum spin x (x = the drum speed for the media that’s being used)
>> plot headtest
>> head install on
>> drum stop
>> drum idle

Note: If plot headtest is not available, use the following procedure:

  1. Stop the drum.
  2. Position the head in front of the plate.
  3. Type
    >> head install on
  4. Record the sum and error values.

To check the correct sum and error values for a Trendsetter-based device, refer to the device installation guide.

For a 300-micron plate on a Lotem Quantum, expect an error of:

  • –130 when the z-stage is forward
  • 0 when the z-stage is back
  • On a Lotem Quantum Spectrum, expect an error value of –50 when the z-stage is forward
Note: On Lotem-based devices, to confirm that you’re using the correct spacer configuration for the media thickness you’re using, contact your PSS.

For more detail, refer to Z-stage spacer size identification on Lotem Quantum

Check the edge detection.
  1. Run the appropriate script:
    • Lotem-based devices: EdgeGrooveQuantum800.spt
    • Trendsetter-based devices: EdgeGroove3244.spt
  2. To interpret the script, refer to: Finding the Correct Drum Position for Edge Detection (Lotem Quantum)

Refer to the script help for a sample graph:

Check the health of the focus laser diode.
  1. Run FocusPI.spt.
  2. Check the Service Shell output window and inspect the graph PeakvBoth.

 

Sample result graphs:

Tune the focus servo and check the focus actuator motion. Run ForcePing.spt.
  1. Inspect the graphs.
    Note: An error message is generated if the script fails.
  2. If the noise in the upper graph (Response LPOS StDev vs. Set LPOS) exceeds 0.5 microns, escalate to the Response Centre.

Check the hvtrim and contrast ratio.
  1. Reset the value of bpd beamposition to 0.
  2. Run ContvHVtrim.spt (contrast vs HVtrim). When the script prompts you for a laser power value:
    • If the customer only uses one media pick the laser power for that media.
    • If the customer runs more than one media, pick a laser power value that is an average for the media that they run.
  3. Look at the 3rd graph (offvHV; U-shaped curve) and select the value at the lowest point (the bottom of the curve).
  4. Apply the new value using the command TH2>set lvd hvtrim xx.x
    Note: This is the voltage that drives the light valve.
  5. Run ContvBP.spt (contrast vs beamposition) to confirm that the beam position with the lowest leakage is 0. Adjest the beam position, if necessary.
  6. Check the media setup (for example, sd, power, slope and curve).

The value of LVDhvtrim should match the value at the low peak on the graph.

Note: This will give the head wider latitude for correcting the stroke, which will make it more reliable.

Sample graph OffvHV

Check the health of the write laser diode.
  1. Adjust the beam steering if necessary (use the command TH2> set bpd beamposition).

If you are using a laptop the results will be automatically stored to the laptop. You will need to copy them over to the workstation.

 

Internal checks

Internal Checks Action Look for Procedure or

Parts

Check the beam steering hold current.

Use the following procedure: Checking the beam steering hold current on a TH2 head

If the incorrect value for the BPD hold current is used, the system will take longer to settle. This may cause a time out.

Correcting the hold current will prevent this.

View a sample of the code readout for these commands.
Check the ECOs/FCOs for the head.

Check that all ECOs/FCOs have been implemented on the head.

A list of FCOs is available on the Thermal 2 FCOs page.

 
Update the ldd calibrate values for the laser startup sequence.
  1. Record the current values:
    Type the command TH2> set ldd and note the value for ldd currentfactor
  2. Run ldd calibrate. See How to Run LDD Calibrate .
  3. Compare these results to the previous values and note any differences.
 

 

Save the NVS settings. In Splot or Service Shell, run: NVSDump.spt    
Check that all parameter values are valid. In Service Shell, run: parmcheck.spt
  1. Ensure that the results from NVSDump.spt are in the results folder.
  2. Run parmcheck.spt.
  3. Check the script results and investigate any parameters listed in yellow or red.
 

Sample Script Graphs

Script name Good Poor/Fail Notes
BBSProfile.spt   bbsprofile_stuck_pixel.png  
BeamBalance.spt BeamBalance_good.pdf If you get results that look like failed_1 or failed_2, escalate to the Response Centre.
ContvHVtrim.spt OffvHV.jpg   The value of LVDhvtrim should match the value at the low peak on the graph.

EdgeMCE.spt

edgemce_good.pdf edgemce_poor.png  

FErrorvSR.spt

FSumvSR_good.pdf  

FocusPI.spt

focuspi_good.png

In this example, the maximum current applied to the focus laser is 47 mA.

focuspi_poor.png

In this example, the maximum current value is more than 117 mA.

To run FocusPI.spt:
  1. Run FocusPI.spt in Splot or Service Shell.
    Note: You do not need to load a plate to run this script.
  2. The script will produce three graphs, PeakvPeak, PeakvAvg and PeakvBoth. Open and inspect PeakvBoth.
  3. You will see two lines that zig-zag from top to bottom. Check that the highest value read at the bottom, current (mA), is not greater than 70 mA. If the current is greater than 70 mA, contact the Response Centre.

LposStatsvLpos.spt

LposStatsvLpos_good.png LposStatsvLpos_poor.png  

    Copyright © 2005 Creo Inc. All rights reserved. Confidential. Back to top.